OpenStudy (anonymous):

square root(15-2x)=X

OpenStudy (anonymous):

To start, we want to square both sides of our equation, to get: \[15-2x=x^2\] From here, we can add 2x to both sides and subtract 15 from both sides to get this in a form that we can work with: \[x^2 +2x -15 = 0\] From here, try solving this quadratic by factoring. I'll get you started: \[(x+?)(x-?)=0\] You're looking for two values that multiply to give -15 and add to give +2.

OpenStudy (whpalmer4):

When you're done, don't forget to test the solutions to the quadratic to see if they also solve the original equation, or if they are extraneous. Need to do this whenever you square both sides of an equation while solving...